# Quantum Cosmology, Decoherence, and the Emergence of Classical Spacetime

### v1

@@ VI Discussion and Outlook
We have shown that classical cosmology emerges dynamically through environment-induced decoherence. The stochastic description of inflationary perturbations arises as the late-time effective theory of a decohered quantum system [10]. Our results clarify the conceptual foundations of quantum cosmology and provide a concrete bridge between universal wavefunctions and classical spacetime histories.
Future directions include extending the analysis to reheating, tensor perturbations, and quantum-information-theoretic characterizations of cosmological decoherence.
VII Conclusion
@@
In this work we have addressed the longstanding question of how classical spacetime and classical cosmological perturbations emerge from quantum cosmology. Starting from standard quantum-cosmological boundary conditions, including the no-boundary and tunneling proposals, we emphasized that semiclassical WKB behavior and squeezing of perturbations are not by themselves sufficient to explain classicality.
@@
We demonstrated that the emergence of classicality is largely insensitive to the details of coarse graining, occurring both for horizon-based and effective-field-theory–motivated splits between system and environment. Classical cosmological histories therefore arise as a generic consequence of inflationary dynamics in the presence of additional degrees of freedom, rather than as a fine-tuned outcome of specific assumptions.
Our analysis clarifies the distinct roles played by quantum-cosmological boundary conditions and decoherence. Boundary conditions determine the relative amplitudes of semiclassical branches in the universal wavefunction, while decoherence determines which of these branches become classical. In particular, expanding cosmological histories are dynamically selected as classical due to their enhanced entanglement with environmental degrees of freedom, providing an emergent arrow of time without imposing one at the fundamental level.
Taken together, these results place the foundations of classical cosmology on the same footing as other emergent phenomena in quantum physics. They provide a coherent and quantitatively controlled framework connecting quantum cosmology, decoherence, and the classical universe described by cosmological observations.

### v2

@@ V Discussion and outlook
We have demonstrated that environment-induced decoherence provides an efficient and robust mechanism for the dynamical emergence of classical cosmology. The stochastic description of inflationary perturbations arises as the late-time effective theory of a decohered quantum system [11, 20]. Our results clarify the conceptual foundations of quantum cosmology and provide a concrete bridge between universal wavefunctions and classical spacetime histories.
Future directions include extending the analysis to the reheating epoch, computing quantum-information-theoretic measures of cosmological decoherence (entanglement entropy, quantum discord) using the branch-overlap framework developed here, and extending the branch-overlap calculation to tensor perturbations — where one expects the same closed form in the massless de Sitter limit, connecting gravitational-wave decoherence to primordial $B$-mode observables via the tensor-to-scalar ratio $r$.
VI Conclusion
@@
In this work we have studied how classical spacetime and classical cosmological perturbations emerge from quantum cosmology. Starting from standard quantum-cosmological boundary conditions, including the no-boundary and tunneling proposals, we emphasized that semiclassical WKB behavior and squeezing of perturbations are not by themselves sufficient to explain classicality.
@@
We demonstrated that the emergence of classicality is largely insensitive to the details of coarse graining, occurring both for horizon-based and EFT-motivated splits between system and environment. Classical cosmological histories therefore arise as a generic consequence of inflationary dynamics in the presence of additional degrees of freedom, rather than as a fine-tuned outcome of specific assumptions.
Our analysis clarifies the distinct roles played by quantum-cosmological boundary conditions and decoherence. Boundary conditions determine the relative amplitudes of semiclassical branches in the universal wavefunction, while decoherence determines when the overlap between these branches becomes negligibly small. The explicit evaluation of the branch-overlap factor $|\mathcal{D}_{k}(z)|$ shows that this overlap vanishes in the superhorizon limit for expanding-branch modes — as a power law $z^{\nu}$ for massive fields and as $z^{1/2}$ exactly for the massless case — while remaining of order unity on the contracting branch. The accumulated functional $\Gamma_{+-}$ crosses the classicality threshold within half an e-fold and grows without bound, establishing the emergent arrow of time as a derived consequence of inflationary squeezing rather than an inference from the formal structure of the equations.
Taken together, these results provide a more explicit framework connecting quantum cosmology, decoherence, and the classical universe described by cosmological observations. In particular, they sharpen how coarse graining and branch decoherence enter the passage from a universal wavefunction to an effective ensemble of classical cosmological histories.
Acknowledgements.
The author thanks Luiz Felipe Demétrio for correspondence regarding bouncing cosmology and Wheeler–DeWitt quantum bounce scenarios, which motivated the discussion of branch decoherence in the context of contracting-phase cosmologies in section IV.
Appendix A Derivation of the decoherence time estimate
For convenience we record the intermediate steps leading from the EFT-based coarse-grained estimate to the explicit decoherence times in table 1. Starting from
$\frac{\mathrm{d}\Gamma_{k_{L}}}{\mathrm{d}\mathcal{N}}\sim\frac{\lambda_{\mathrm{eff}}^{2}\Lambda_{\mathrm{phys}}}{H^{5}}\,|\Delta\zeta_{k_{L}}|^{2}\,a^{4}(\mathcal{N}),$ (56)
and writing $a(\mathcal{N})=a_{\ast}\,\mathrm{e}^{\Delta\mathcal{N}}$ relative to horizon exit of the long mode, one finds
$\frac{\mathrm{d}\Gamma_{k_{L}}}{\mathrm{d}\mathcal{N}}\sim\frac{\lambda_{\mathrm{eff}}^{2}\Lambda_{\mathrm{phys}}}{H^{5}}\,|\Delta\zeta_{k_{L}}|^{2}\,a_{\ast}^{4}\,\mathrm{e}^{4\Delta\mathcal{N}}.$ (57)
Using horizon exit, $a_{\ast}H=k_{L}$, as the reference scale and absorbing order-unity factors into the normalization,
$\frac{\mathrm{d}\Gamma_{k_{L}}}{\mathrm{d}\mathcal{N}}\sim\hat{\lambda}^{2}\,\frac{\Lambda_{\mathrm{phys}}}{H}\,|\Delta\zeta_{k_{L}}|^{2}\,\mathrm{e}^{4\Delta\mathcal{N}},\quad\hat{\lambda}\equiv\frac{\lambda_{\mathrm{eff}}}{H^{2}}.$ (58)
Integrating from $\Delta\mathcal{N}^{\prime}=0$ to $\Delta\mathcal{N}$ yields
$\Gamma_{k_{L}}(\Delta\mathcal{N})=\frac{\hat{\lambda}^{2}}{4}\,\frac{\Lambda_{\mathrm{phys}}}{H}\,|\Delta\zeta_{k_{L}}|^{2}\,\left(\mathrm{e}^{4\Delta\mathcal{N}}-1\right),$ (59)
which is the expression used in equation (39). Solving for $\Delta\mathcal{N}_{\mathrm{dec}}$ subject to $\Gamma_{k_{L}}(\Delta\mathcal{N}_{\mathrm{dec}})\simeq 1$ reproduces equation (39). Using $\zeta_{\mathrm{rms}}\simeq\sqrt{A_{s}}\simeq 5\times 10^{-5}$ and $\hat{\lambda}=\tfrac{3}{2}m_{\sigma}^{2}/H^{2}$ yields the numerical values in table 1 and the mass-based examples in section III.5.